OpenStudy (barreraa):

For the function f(x) = 2 – 3x, find f(4). (Points : 4) –10 –5 12 14

OpenStudy (barreraa):

@BlossomCake

OpenStudy (diamondboy):

wherever u see x plug in 4

OpenStudy (agl202):

f(4) = 2 - 3(4) f(4) = 2 - 12 f(4) = -10 Answer (A)

OpenStudy (anonymous):

Okay, so for this one we have to substitute f(4) for all the 'x' in the equation and solve.

OpenStudy (anonymous):

Please do not give out a direct answer. @Agl202